1. George and Ringo are stranded on a deserted island where they survive by picking mangoes and catching fish. Ringo can get 4 fish or 8 mangoes in a day. George can get 10 fish or 10 mangoes in a day. When working separately, they each choose to spend half of their day getting fish and half getting mangoes.

a. How many Fish and mangoes does each get when working alone?

b. Give the marginal cost of fish for both George and Ringo.

c. Who has a comparative advantage in fish and who in mangoes?

d. If George and Ringo coordinate their efforts, draw the production possibility frontier for the two of them combined. Label all slopes and intercepts on your graph.
Answer #1

a) When George and Ringo works alone then they spend half time to pick mangoes and other half catch fish.

Ringo: Fish = 2 and Mangoes = 4

George: Fish = 5 and Mangoes = 5

b) George;

10 fishes = 10 mangoes

1 fish = 10/10 = 1 mango

Marginal cost of 1 fish is 1 mango.

Ringo;

4 fish = 8 mangoes

1 fish = 8/4 = 2 mangoes

Marginal cost of 1 fish is 2 mangoes.

c) A country has a comparative advantage in producing that good if the opportunity cost of producing that good is lower in that country as compared to another country.

George has comparative advantage in the production of fish because of lower opportunity cost or marginal cost of producing 1 fish. Ringo has comparative advantage in the production of mangoes.

d)

Mangoes 18 li카카 0 14 Fishes

Maximum production of fishes = 4 fishes by Ringo + 10 fishes by George = 14 fishes

Maximum production of mangoes = 8 mangoes by Ringo + 10 mangoes by George = 18 mangoes

slope = - 14/18 = - 7/9